# Comparing the latent features of universal machine-learning interatomic potentials

### v1

@@ Abstract
The past few years have seen the development of “universal” machine-learning interatomic potentials (uMLIPs) capable of approximating the ground-state potential energy surface across a wide range of chemical structures and compositions with reasonable accuracy. While these models differ in the architecture and the dataset used, they share the ability to compress a staggering amount of chemical information into descriptive latent features. Herein, we systematically analyze what the different uMLIPs have learned by quantitatively assessing the relative information content of their latent features with feature reconstruction errors as metrics, and observing how the trends are affected by the choice of training set and training protocol. We find that the uMLIPs encode chemical space in significantly distinct ways, with substantial cross-model feature reconstruction errors. When variants of the same model architecture are considered, trends become dependent on the dataset, target, and training protocol of choice. We also observe that fine-tuning of a uMLIP retains a strong pre-training bias in the latent features. Finally, we discuss how atom-level features, which are directly output by MLIPs, can be compressed into global structure-level features via concatenation of progressive cumulants, each adding significantly new information about the variability across the atomic environments within a given system.
@@ I Introduction
Machine learning (ML) has brought transformative changes to the field of computational materials science. Data-driven algorithms enable simulations and electronic structure calculations with near ab initio accuracy to reach unprecedented length and time scales Jacobs2025; Schmidt2019MLinMaterials; Qu2024; Mahmoud2020; How2025; shnorb; Suman2025. A particularly active frontier in this field is the development of “universal” machine-learned interatomic potentials (uMLIPs) that aim for applicability across the entire chemical composition space DPA3; UMA; MACE; mace_multihead; gnome; PET-MAD-2025; MatterSim; OpenLAMv1. Such models alleviate the need for laborious dataset construction and model training from scratch, and instead allow zero-shot deployment or fine-tuning of existing uMLIPs, further accelerating applications across diverse areas of chemistry and materials science.
@@
To date, several dozens of uMLIPs have been developed, each employing distinct atomic descriptors, neural network (NN) architectures, and training datasets of varying scope and composition MPTRAJ; Alexandria; MAD; MATPES; SPICE; OC20; ODAC; omat24; OMC25; OMOL. Despite their rapidly growing diversity, most uMLIPs have been evaluated primarily on curated benchmarking suites that assess their accuracy across standard material modeling tasks matbench; LAMBench; chiang2025mliparena. This limited perspective motivates a deeper examination of the similarities and differences among uMLIPs, particularly how they represent and interpret the chemical space.
@@
Regardless of their architectural details, all uMLIPs share a key characteristic: they encode the multidimensional description of atomic systems, comprised of types and positions of constituent atoms, into a compact latent representation containing only a few hundred features. These latent representations embody the distinct ways in which different uMLIPs perceive and organize the chemical space, and the utility of these latent features has been demonstrated in various contexts: uncertainty quantification llpr; fast_uncertainty; uncertainty_nn; kulichenko2024uncertainty; Janet2019ChemSci; Feldmann2025, outlier detection Vazquez-Salazar2025; Schwalbe_Koda_2025, active learning Xie2021BayesianFF; Vandermause2020; kulichenko2024uncertainty, dataset curation kulichenko2024data; zou2025data; Glielmo2022, and system classification de+16jci; stability_classification.
@@
In this work, we leverage the latent features of uMLIPs to systematically and quantitatively analyze the similarities and differences in how these models perceive and structure the chemical space. We adopt the global and local feature reconstruction errors of Goscinski et al. Goscinski_2021 as the primary metrics to assess the information content of the latent features of four uMLIPs: MACE-MP-03b MACE, PET-MAD PET-MAD-2025, DPA-3.1 DPA3, and UMA-S-1P1 UMA. We investigate the mutual reconstructability of the uMLIP latent features, as well as those of model variants trained with different datasets and training strategies. We also extend the evaluation to the latent features of PET-MAD and those of PET-MAD-DOS PET-MAD-DOS-2025, a model with nearly identical architecture and training dataset, but trained for a different target property. We then study the evolution of the latent space during fine-tuning, and explore how the atomic latent features can be efficiently aggregated into global features from the information content standpoint.
@@
We find that all uMLIPs uniquely encode chemical space with significant cross-model reconstruction errors, where some models are more accurately linearly reconstructed by others. When considering the variants of model architectures targeting different datasets, we observe that single-task models (MACE-mp-03b) and multi-task models (MACE-mh-1 mace_multihead, DPA-3.1) share relatively consistent representations across datasets, while the mixture of linear experts (MoLE) model (UMA-S-1P1) specializes the features far more distinctly. We observe that the information content of “last-layer” and “backbone” features in a model architecture is similar, yet the backbone features encode more information that last-layer features cannot reconstruct. We also demonstrate that fine-tuning of the uMLIP exhibits a strong pre-training bias in the latent feature space. Finally, we demonstrate the compression of local, atomic features into global structure-level features by concatenating their progressive cumulants, resulting in an information-rich structural descriptor that obeys desirable scaling laws and preserves the knowledge of configurational inhomogeneity.
@@ II Methods
Table 1: Key characteristics of uMLIPs considered in this work. Dataset sizes are reported in the number of structures. For UMA-S-1P1, 6M denotes the active parameters used per structure during inference, while 150M is the total across all experts.
MACE-MP-03b PET-MAD DPA-3.1 UMA-S-1P1
@@
Dataset(s) MPtrj (1.58M) MPTRAJ MAD (0.096M) MAD OpenLAM-v1 (163M) OpenLAMv1 OMat24, OC20, ODAC25, OMC25, OMol25 (total 569M) omat24; OC20; ODAC; OMC25; OMOL
Equivariance E(3)-equivariant Rotationally unconstrained; equivariance learned by data augmentation Invariant under rotation and translation E(3)-equivariant
@@
For our analysis, we mainly consider MACE-MP-03b, PET-MAD, DPA-3.1, and UMA-S-1P1, which constitute a non-exhaustive yet diverse set of uMLIPs that differ vastly in their message-passing NN architectures, learning strategies, and training datasets (see Table 1). MACE-MP-03b MACE is built on an E(3)-equivariant message-passing NN that constructs higher body-order correlations using equivariant tensor products and body-ordered polynomials. PET-MAD PET-MAD-2025 is based on a rotationally unconstrained transformer-based graph neural network (GNN), where strict geometrical equivariance is not enforced by architecture but is effectively learned through rotational data augmentation during training PET. DPA-3.1 DPA3 is an attention-based model of the Deep Potential (DP) family DeepMD-kit-v3 that produces fully rotation- and translation-invariant atomic descriptors via a gated attention mechanism. UMA-S-1P1 UMA is implemented upon a large-scale E(3)-equivariant mixture-of-experts architecture built on the eSEN eSEN backbone that uses a MoLE design. For UMA-S-1P1 and DPA-3.1, which are a MoLE model and a multi-head model, respectively, we use the expert or the head corresponding to the OMat24 dataset omat24, given that MACE-MP-03b and PET-MAD have been trained on similar materials-oriented datasets.
We define the latent features of each uMLIP as the tokens extracted from a given layer of the model architecture. These features have been processed through multiple layers of message-passing to encode the information aggregated from neighboring atoms within the defined cutoff radius. We primarily focus on the last-layer features, corresponding to the representation most refined for the regression task obtained after all message-passing iterations and subsequent multilayer perceptron (MLP) transformations. Since all considered uMLIPs adopt a locality ansatz, we extract atomic features $\boldsymbol{\xi}(A_{i})^{F}$ for each atom $i$ in structure $A$ via a forward-pass through the model $F$. To enable a consistent, scale-invariant comparison across the models, the features are centered to zero mean and globally scaled so that their total variance across all dimensions equals one on the analysis dataset for each uMLIP, thereby preserving the relative scales between feature dimensions.
@@ II.1 Feature space comparison metrics
In our analysis, we primarily take a quantitative approach in studying the similarities and differences between the latent space of uMLIPs. To this end, we compute the global feature reconstruction error (GFRE) and the local feature reconstruction error (LFRE) proposed by Goscinski et al. Goscinski_2021. These metrics assess the relative expressive power of feature spaces by determining how well one set of features can reconstruct another.
@@
To restrict the learning problem to modeling only relative displacements on the local manifold patch, each neighborhood in centered by subtracting its mean:
@@
We clarify that the definition of “global” and “local” in the context of feature reconstruction errors does not correspond to the usual atomic versus system-wide considerations in atomistic ML (see SI V for a schematic overview). Rather, “global” denotes a single linear transformation applied across the entire dataset of atomic features, whereas “local” involves fitting separate linear mappings for localized neighborhoods within the high-dimensional feature space. Both metrics can operate on both atomic-level and structure-level features, but differ in how they assess reconstructability across the feature manifold.
Owing to the feature normalization, both the GFRE and LFRE are usually bounded between 0 and $1$, with 0 signifying perfect linear reconstruction, thus comparable information between the feature spaces. However, the upper bounds can be exceeded in case of reconstruction failure due to ill-conditioned linear mappings, which indicates large dissimilarity or incompatibility between the feature spaces. For our analysis, we use the GFRE and LFRE implementations in scikit-matter scikit-matter with ridge regression and default 2-fold cross-validation to determine the ridge regularization, with regularization $\alpha$ values scanned over a logarithmic range from $10^{-9}$ to 1.
@@ II.2 Feature space projection
In some examples, we complement this quantitative analysis by a visual mapping of the uMLIP features to a lower-dimensional space using Principal Covariates Regression (PCovR) pcovr; cersonsky_pcovr. This dimensionality reduction technique interpolates between two objective functions: maximizing variance retention in the original features via principal component analysis (PCA), and optimizing linear regression performance on a target property. In our context, by mixing in the regression term in addition to the PCA term, we gently enforce a degree of consistency across the maps produced for different models to aid interpretation.
@@
where $\widehat{X}=\mathbf{T}P_{X}^{\top}$ is the approximation of the original feature space $X$ (PCA term) and $\widehat{Y}=\mathbf{T}P_{Y}^{\top}$ is the prediction of targets (linear regression term). The interpolation between the two terms is controlled by a tunable parameter $\alpha\in\left[0,1\right]$. In our analysis, we use scikit-matter implementation with $\alpha=0.5$ and $k=2$ principal components.
To prepare the inputs, atomic features are aggregated to the structure-level by computing the cumulants introduced in Sec. III.6 across atoms within each structure. The target property $Y$ is defined as the cohesive energies per atom, calculated by subtracting an atomic baseline energy from the total potential energy of each structure and normalizing by the number of atoms.
@@ III.1 Latent feature comparison between uMLIPs
Figure 1: Heat maps of GFRE (left) and LFRE (right) for atomic last-layer latent features of MACE-MP-03b, PET-MAD, DPA-3.1, and UMA-S-1P1, computed with the atomic environments on the test split of the MAD dataset. Each cell represents the reconstruction error when mapping latent features of the “source” (row) to “target” (column).
@@
We first quantitatively compare the latent features of different uMLIPs by computing the GFRE and LFRE between the four models considered. The errors are computed for structures in the test subset of the Massive Atomic Diversity (MAD) dataset MAD, which contains chemical systems with a broad coverage across 85 elements and exhibits vast structural diversity. We exclude the single-atom structures (incompatible with UMA) and polonium- or radon-containing structures (not supported by MACE trained on MPtrj) to yield 258,561 atomic environments from 9,461 structures. We then perform random stratified sampling across the subsets of MAD dataset at the atomic level to select 1,000 atomic environments for $\mathcal{D}_{\text{test}}$. All remaining environments are used as $\mathcal{D}_{\text{train}}$ to compute the regression weights $\widehat{P}_{FF^{\prime}}$.
The heat maps in Fig. 1 reveal distinct trends in the feature reconstruction performances across the uMLIPs. Overall, the values are large (average of 0.66 for GFRE and 0.37 for LFRE off-diagonal cross-terms), indicating that each model encodes the chemical space in a unique manner. The LFREs are generally lower than the GFREs, which reflects the ability of the LFRE to implicitly account for nonlinear relationships between the feature spaces. The lowest GFREs and LFREs are observed when MACE-MP-03b features are used as targets, suggesting that the other uMLIPs can reliably recover the information encoded in its features. Conversely, the highest errors are observed when DPA-3.1 features serve as targets.
@@
The average GFREs across the off-diagonal targets for each source model are 0.70 for MACE-MP-03b, 0.60 for PET-MAD, 0.67 for DPA-3.1, and 0.68 for UMA-S-1P1. For the LFRE, they are 0.35, 0.33, 0.38, and 0.44, respectively. PET-MAD exhibits marginally lower GFREs and LFREs in reconstructing the features of all other uMLIPs, despite being trained on the smallest dataset among the models considered. To rule out the fact that the low PET-MAD reconstruction errors are lower only because we are using the MAD dataset as a benchmark, we repeat the analysis using the Alexandria subset Alexandria (cf. SI VI). We observe similar trends, indicating that the smaller MAD dataset can still sufficiently represent the configuration space of much larger datasets, and that its choice for our analysis does not introduce a significant bias.

### v2

@@ Abstract
The past few years have seen the development of “universal” machine-learning interatomic potentials (uMLIPs) capable of approximating the ground-state potential energy surface across a wide range of chemical structures and compositions with reasonable accuracy. While these models differ in the architecture and the dataset used, they share the ability to compress a staggering amount of chemical information into descriptive latent features. Herein, we systematically analyze what the different uMLIPs have learned by quantitatively assessing the relative information content of their latent features with feature reconstruction errors, and observing how the trends are affected by the choice of training set and training protocol. We find that uMLIPs encode the chemical space in significantly distinct ways, with substantial cross-model feature reconstruction errors. When variants of the same model architecture are considered, trends become dependent on the dataset, target, and training protocol of choice. We also observe that fine-tuning of a uMLIP retains a strong pre-training bias in the latent features. Finally, we discuss how atom-level features, which are directly output by MLIPs, can be compressed into global structure-level features via concatenation of progressive cumulants, each adding significantly new information about the variability across the atomic environments within a given system.
@@ I Introduction
Machine learning (ML) has brought transformative changes to the field of computational materials science. Data-driven algorithms enable simulations and electronic structure calculations with near ab initio accuracy to reach unprecedented length and time scales Jacobs et al. (2025); Schmidt et al. (2019); Qu and Krishnapriyan (2024); Ben Mahmoud et al. (2020); How et al. (2025a); Schütt et al. (2019); Suman et al. (2025). A particularly active frontier in this field is the development of “universal” machine-learned interatomic potentials (uMLIPs) that aim for applicability across the entire chemical composition space Zhang et al. (2025); Wood et al. (2025); Batatia et al. (2025a, b); Merchant et al. (2023); Mazitov et al. (2025a); Yang et al. (2024); Peng et al. (2025b). Such models alleviate the need for laborious dataset construction and model training from scratch, and instead allow zero-shot deployment or fine-tuning of existing uMLIPs, further accelerating applications across diverse areas of chemistry and materials science.
@@
To date, several dozens of uMLIPs have been developed, each employing distinct atomic descriptors, neural network (NN) architectures, and training datasets of varying scope and composition Deng et al. (2023); Schmidt et al. (2024); Mazitov et al. (2025b); Kaplan et al. (2025); Eastman et al. (2023); Chanussot et al. (2021); Sriram et al. (2023); Barroso-Luque et al. (2024); Gharakhanyan et al. (2025); Levine et al. (2025). Despite their rapidly growing diversity, most uMLIPs have been evaluated primarily on curated benchmarking suites that assess their accuracy across standard material modeling tasks Riebesell et al. (2025); Peng et al. (2025a); Chiang et al. (2025). This limited perspective motivates a deeper examination of the similarities and differences among uMLIPs, particularly how they represent and interpret the chemical space Bonneau et al. (2025).
@@
Regardless of their architectural details, all uMLIPs share a key characteristic: they encode the multidimensional description of atomic systems, comprised of types and positions of constituent atoms, into a compact latent representation containing only a few hundred features. These latent representations embody the distinct ways in which different uMLIPs perceive and organize the chemical space, and the utility of these latent features has been demonstrated in various contexts: uncertainty quantification Bigi et al. (2024); Zhu et al. (2023); Vazquez-Salazar et al. (2022); Kulichenko et al. (2023); Janet et al. (2019); Feldmann et al. (2025), outlier detection Vazquez-Salazar et al. (2025); Schwalbe-Koda et al. (2025), active learning Xie et al. (2021); Vandermause et al. (2020); Kulichenko et al. (2023), dataset curation Kulichenko et al. (2024); Zou and Marzouk (2025); Glielmo et al. (2022), and system classification De et al. (2017); Bartel et al. (2020).
@@
In this work, we leverage the latent features of uMLIPs to systematically and quantitatively analyze the similarities and differences in how these models perceive and structure the chemical space. While recent complementary works have focused on discerning whether there exists a converged or aligned representation of matter across the uMLIPs Edamadaka et al. (2025); Li and Walsh (2025), here we adopt a purely statistical approach, using the global and local feature reconstruction errors of Goscinski et al. Goscinski et al. (2021) to assess the information content across the latent features of four uMLIPs: MACE-MP-0b3 Batatia et al. (2025a), PET-MAD Mazitov et al. (2025a), DPA-3.1 Zhang et al. (2025), and UMA-S-1P1 Wood et al. (2025). We investigate the mutual reconstructability of the uMLIP latent features, as well as those of model variants trained with different datasets and training strategies. We also extend the evaluation to the latent features of PET-MAD and those of PET-MAD-DOS How et al. (2025b), a model with nearly identical architecture and dataset but trained for a different target property. We then study the evolution of the latent space during fine-tuning, and explore how the atomic latent features can be efficiently aggregated into global features from the information content standpoint.
@@
We find that all uMLIPs uniquely encode chemical space with significant cross-model reconstruction errors, where some models are more accurately linearly reconstructed by others. When considering the variants of model architectures targeting different datasets, we observe that single-task models (MACE-MP-0b3) and multi-task models (MACE-mh-1 Batatia et al. (2025b), DPA-3.1) share relatively consistent representations across datasets, while the mixture of linear experts (MoLE) model (UMA-S-1P1) specializes the features far more distinctly. We observe that the information content of “last-layer” and “backbone” features in a model architecture is similar, yet the backbone features encode more information that last-layer features cannot reconstruct. We also demonstrate that fine-tuning of the uMLIP exhibits a strong pre-training bias in the latent feature space. Finally, we demonstrate the compression of local, atom-level features into global, structure-level features by concatenating their progressive cumulants, resulting in an information-rich structural descriptor that obeys desirable scaling laws and preserves the knowledge of configurational inhomogeneity.
@@ II Methods
Table 1: Key characteristics of the uMLIPs considered in this work. Dataset sizes are reported by the number of structures. For UMA-S-1P1, 6M denotes the active parameters used per structure during inference, while 150M is the total across all the experts.
MACE-MP-0b3 PET-MAD DPA-3.1 UMA-S-1P1
@@
Dataset(s) MPtrj (1.58M) Deng et al. (2023) MAD (0.096M) Mazitov et al. (2025b) OpenLAM-v1 (163M) Peng et al. (2025b) OMat24, OC20, ODAC25, OMC25, OMol25 (total 569M) Barroso-Luque et al. (2024); Chanussot et al. (2021); Sriram et al. (2023); Gharakhanyan et al. (2025); Levine et al. (2025)
Equivariance E(3)-equivariant Rotationally unconstrained; equivariance learned by data augmentation SE(3)-invariant E(3)-equivariant
@@
For our analysis, we mainly consider MACE-MP-0b3, PET-MAD, DPA-3.1, and UMA-S-1P1, which constitute a non-exhaustive yet diverse set of uMLIPs that differ vastly in their message-passing NN architectures, learning strategies, and training datasets (see Table 1). MACE-MP-0b3 Batatia et al. (2025a) is built on an E(3)-equivariant message-passing NN that constructs higher body-order correlations using equivariant tensor products and body-ordered polynomials. PET-MAD Mazitov et al. (2025a) is based on a rotationally unconstrained transformer-based graph neural network (GNN), where strict geometrical equivariance is not enforced by architecture but is effectively learned through rotational data augmentation during training Pozdnyakov and Ceriotti (2023). DPA-3.1 Zhang et al. (2025) is an attention-based model of the Deep Potential (DP) family Zeng et al. (2025) that produces fully rotation- and translation-invariant atomic descriptors via a gated attention mechanism. UMA-S-1P1 Wood et al. (2025) is implemented upon a large-scale E(3)-equivariant mixture-of-experts architecture built on the eSEN Fu et al. (2025) backbone that uses a MoLE design. For UMA-S-1P1 and DPA-3.1, which are a MoLE model and a multi-head model, respectively, we use the expert or the head corresponding to the OMat24 dataset Barroso-Luque et al. (2024), given that MACE-MP-0b3 and PET-MAD have been trained on similar materials-oriented datasets.
We define the latent features of each uMLIP as the tokens extracted from a given layer of the model architecture. These features have been processed through multiple layers of message-passing to encode the information aggregated from neighboring atoms within the defined cutoff radius. We primarily focus on the last-layer features, corresponding to the representation most refined for the regression task obtained after all message-passing iterations and subsequent multilayer perceptron (MLP) transformations. Since all considered uMLIPs adopt a locality ansatz, we extract atomic features $\boldsymbol{\xi}(A_{i})^{F}$ for each atom $i$ in structure $A$ via a forward-pass through the model $F$.
To enable a consistent, scale-invariant comparison across the models, the features are centered to zero mean and globally scaled so that their total variance across all dimensions equals one on the analysis dataset for each uMLIP, thereby preserving the relative scales between feature dimensions.
@@ II.1 Feature space comparison metrics
In our analysis, we primarily take a quantitative approach in studying the similarities and differences between the latent space of uMLIPs. To this end, we compute the global feature reconstruction error (GFRE) and the local feature reconstruction error (LFRE) proposed by Goscinski et al. Goscinski et al. (2021). These metrics assess the relative expressive power of feature spaces by determining how well one set of features can reconstruct another.
@@
To restrict the learning problem to modeling only relative displacements on the local manifold patch, each neighborhood is centered by subtracting its mean:
@@
We clarify that the definitions of “global” and “local” in the context of feature reconstruction errors do not correspond to the usual atomic versus system-wide considerations in atomistic ML (see SI V for a schematic overview). Rather, “global” denotes a single linear transformation applied across the entire dataset of atomic features, whereas “local” involves fitting separate linear mappings for localized neighborhoods within the high-dimensional feature space. Both metrics can operate on both atomic-level and structure-level features, but differ in how they assess reconstructability across the feature manifold.
Owing to the feature normalization, both the GFRE and LFRE are usually bounded between 0 and $1$, with 0 signifying perfect linear reconstruction, thus comparable information between the feature spaces. However, the upper bounds can be exceeded in case of reconstruction failure due to ill-conditioned linear mappings, which indicates large dissimilarity or incompatibility between the feature spaces. For our analysis, we use the GFRE and LFRE implementations in scikit-matter Goscinski et al. (2023) with ridge regression and default 2-fold cross-validation to determine the ridge regularization, with regularization $\alpha$ values scanned over a logarithmic range from $10^{-9}$ to 1, and the final value was selected as the smallest $\alpha$ that provided numerical stability.
@@ II.2 Feature space projection
In a few examples, we complement this quantitative analysis by a visual mapping of the uMLIP features to a lower-dimensional space using Principal Covariates Regression (PCovR) de Jong and Kiers (1992); Cersonsky et al. (2021). This dimensionality reduction technique interpolates between two objective functions: maximizing variance retention in the original features via principal component analysis (PCA), and optimizing linear regression performance on a target property. In our context, by mixing in the regression term in addition to the PCA term, we gently enforce a degree of consistency across the maps produced for different models to aid interpretation.
@@
where $\widehat{X}=\mathbf{T}P_{X}^{\top}$ is the approximation of the original feature space $X$ (PCA term) and $\widehat{Y}=\mathbf{T}P_{Y}^{\top}$ is the prediction of targets (linear regression term). The matrices $P_{X}$ and $P_{Y}$ represent projection weights that map latent coordinates back to the original feature and target dimensions. The interpolation between the two terms is controlled by a tunable parameter $\alpha\in\left[0,1\right]$. In our analysis, we use scikit-matter implementation with $\alpha=0.5$ and $k=2$ principal components.
To prepare the inputs, atomic features are aggregated to the structure-level by computing the cumulants introduced in Sec. III.6 across atoms within each structure. The target property $Y$ is defined as the cohesive energies per atom, calculated by subtracting an atomic baseline energy from the reference total potential energy of each structure and normalizing by the number of atoms.
@@ III.1 Latent feature comparison between uMLIPs
Figure 1: Heat maps of the global feature reconstruction error (GFRE, left) and local feature reconstruction error (LFRE, right) for the atomic last-layer latent features of MACE-MP-0b3, PET-MAD, DPA-3.1, and UMA-S-1P1, computed for the atomic environments from the test subset of the MAD dataset. Each cell represents the reconstruction error when mapping latent features of the “source” (row) to “target” (column).
@@
We first quantitatively compare the latent features of different uMLIPs by computing the GFRE and LFRE between the four models considered. The errors are computed for structures in the test subset of the Massive Atomic Diversity (MAD) dataset Mazitov et al. (2025b), which contains chemical systems with a broad coverage across 85 elements and exhibits vast structural diversity. We exclude the single-atom structures (incompatible with UMA) and polonium- or radon-containing structures (not supported by MACE trained on MPtrj) to yield 258,561 atomic environments from 9,461 structures. We then perform random stratified sampling across the subsets of MAD dataset at the atomic level to select 1,000 atomic environments for $\mathcal{D}_{\text{test}}$. All remaining environments are used as $\mathcal{D}_{\text{train}}$ to compute the regression weights $\widehat{P}_{FF^{\prime}}$.
The heat maps in Fig. 1 reveal distinct trends in the feature reconstruction performances across the uMLIPs. Overall, the values are large (average of 0.66 for GFRE and 0.37 for LFRE off-diagonal cross-terms), indicating that each model encodes the chemical space in a unique manner. That latent spaces of different models differ significantly is not surprising. Latent features are only weakly constrained by the learning target: in principle, a model could learn the energy as the first feature, and have all others reflect completely unrelated components of the molecular structure. Therefore, the latent space reflects the priors explicitly and implicitly encoded in the model architecture, the choice of hyperparameters, and the training strategy.
@@
The LFREs are generally lower than the GFREs, which reflects the ability of the LFRE to implicitly account for nonlinear relationships between the feature spaces. The lowest GFREs and LFREs are observed when MACE-MP-0b3 features are used as targets, suggesting that the other uMLIPs can reliably recover the information encoded in its features. Conversely, the highest errors are observed when DPA-3.1 features serve as targets. The average GFREs across the off-diagonal targets for each source model are 0.70 for MACE-MP-0b3, 0.60 for PET-MAD, 0.67 for DPA-3.1, and 0.68 for UMA-S-1P1. For the LFRE, they are 0.35, 0.33, 0.38, and 0.44, respectively. PET-MAD exhibits marginally lower GFREs and LFREs in reconstructing the features of all other uMLIPs, despite being trained on the smallest dataset among the models considered. To rule out the fact that the low PET-MAD reconstruction errors are lower only because we are using the MAD dataset as a benchmark, we repeat the analysis using the Alexandria subset Schmidt et al. (2024) (cf. SI VI). We observe similar trends, indicating that the smaller MAD dataset can still sufficiently represent the configuration space of much larger datasets, and that its choice for our analysis does not introduce a significant bias.
In SI VII, we also present the feature reconstruction errors (FREs) across the models for the backbone features (representations obtained immediately after message-passing and before the MLP readout), which show smaller reconstruction error (average off-diagonal GFRE is 0.54, LFRE is 0.30) and hence relatively higher generalizability across the models.